# Human or LLM as Standardized Patients? A Comparative Study for Medical Education

### v1

@@ 3.3 Evaluation Protocol
4 Experiment Protocol
After validating our virtual SP on SPBench, we conducted a user study to evaluate its educational impact in a real-world setting. This section describes the four-week randomized crossover trial with undergraduate medical students, designed to test whether EasyMED outperforms traditional human SP training.
Study Period Timeline Group A Group B
Baseline Week 0 Pre-Test Pre-Test
Period 1 Weeks 1-2 Intervention A: Intervention B:
EasyMED Training Human SP Training
End of Week 2 Mid-Test Mid-Test
Period 2 Weeks 3-4 Intervention B: Intervention A:
Human SP Training EasyMED Training
End of Week 4 Final Test & Questionnaire
Table 7: Design of the Experiment, illustrating the sequence of interventions (EasyMED Training and Human SP Training) and assessments for Group A and Group B.
4.1 Experimental Design
We ran a four-week randomized crossover experiment comparing EasyMED with traditional human SP training. Each participant experienced both training methods in different phases, enabling within-subject comparisons and controlling for individual differences in baseline competency. The procedure is shown in Table 7.
Patient Cases
During the intervention, students practiced with 20 standardized clinical cases covering core diseases at the scope and difficulty of the National Medical Licensing Examination (see Appendix D).
4.2 Participants
@@
We recruited 20 fourth- and fifth-year undergraduate medical students, all of whom completed a pre-test to assess baseline knowledge and received a 10-minute EasyMED orientation. Students with scheduling conflicts or extreme pre-test scores were excluded, resulting in a final sample of 14 (7 males, 7 females; mean age 23, range 21–-24). All had completed core clinical coursework but had not yet taken the National Medical Licensing Examination.
To ensure baseline equivalence, participants were rank-ordered by pre-test scores and alternately assigned to two sequence groups. Three experienced SP professionals delivered the human SP sessions. All participants received hourly compensation in accordance with ethical regulations.
5 Results and Analysis
This section presents and analyzes the results from Section 4, focusing on whether the Virtual SP-based EasyMED can deliver learning outcomes equivalent to or better than traditional human SP training.
5.1 Evaluating Overall Improvement via OSCE
Figure 4: Boxplots comparing the baseline scores of Group A and Group B. Each circle represents the score of an individual participant. The similarity in the distributions demonstrates that the two groups were comparable at the start of the experiment.
Both groups started from an equivalent baseline We first confirmed that the two groups were comparable at baseline. Baseline equivalence. Figure 4 shows that the two groups had similar initial score distributions. An independent-samples t-test confirmed no significant difference between Group A (mean = 70.56) and Group B (mean = 69.84), t(12) = 0.16, p = 0.88, indicating comparable baselines.
Overall Performance Improvement:
Both training methods produced clear, comparable gains. Specifically, As shown in Table 6, Group A’s average score rose from 70.56 to 87.44 (+16.89), while Group B’s increased from 69.84 to 85.20 (+15.36). Figure 5 illustrates this upward trend across sessions.
@@ 5.1 Evaluating Overall Improvement via OSCE
Improvement at Each Stage
Most gains occurred in the first two weeks, followed by a slowdown in Weeks 3-–4. Table 6 shows that Group A improved by 15.51 points and Group B by 12.17 points in the first two weeks. In Weeks 3–4 the rate of improvement slowed markedly, indicating diminishing returns and underscoring the strong initial impact of both training methods.
Figure 7: This chart shows the total score gain for each group, broken down by training phase. The total height of each bar represents the average total gain for that group. The stacked segments show the portion of the gain that occurred during Phase 1 (Weeks 1-2) and Phase 2 (Weeks 3-4).
Improvement by Skill Level
EasyMED delivered the largest benefits for low-baseline students early on. We split participants into high-baseline (top three) and low-baseline (bottom four) subgroups based on pre-test scores. As shown in Figure 6, low-baseline students using EasyMED gained an average of 21.83 points versus 16.58 in the human SP group. High-baseline students improved more modestly—7.10 for EasyMED versus 6.30 for human SP—showing that EasyMED is particularly advantageous for learners starting at a lower skill level.
Improvement by Training Method
Across the crossover design, EasyMED produced larger score gains than traditional SP training. Although the final post-test scores for Group A (M = 87.44) and Group B (M = 85.20) were not significantly different ($p>.05$), the crossover design allowed us to analyze the score gains from each training method independently (Figure 7). Training with EasyMED led to an average gain of 9.35 points, compared with 6.77 points for training with a human SP. This indicates that, while both methods are effective, EasyMED yields greater skill improvement than traditional SP training.
5.2 Behavioral Analysis via Survey and Logs
To investigate why EasyMED is effective, we analyzed both behavioral logs and questionnaire responses.
Training Flexibility
Students rated EasyMED more usable and used it far more flexibly, with practice extending into evenings and weekends. Students gave EasyMED significantly higher usability scores than human SP (4.8 vs. 3.5, $p<.01$). This preference was reflected in activity patterns (Figure 8): human SP sessions clustered during weekday working hours, whereas EasyMED practice was spread throughout the day and peaked on evenings and weekends. These patterns indicate that EasyMED overcomes the scheduling limits of traditional training, offering more flexible and self-paced learning.
Psychological Safety and Engagement
EasyMED provided a lower-anxiety environment that encouraged deeper engagement and more questioning. Students reported significantly less learning anxiety with EasyMED than with human SP (mean score 2.5 vs. 3.2, $p<.01$). Interviews echoed this, with participants noting they were “not afraid to make mistakes” and felt comfortable “asking follow-up questions repeatedly.” This sense of safety translated into higher engagement: sessions with EasyMED averaged more dialogue turns (54 vs. 47) and longer durations (28:49 vs. 15:17) than human SP sessions (Table 8). Although part of the extra time reflects typing, the higher number of dialogue turns more strongly indicates engagement, suggesting that the low-anxiety environment promotes deeper reasoning and inquiry.
@@ 5.2 Behavioral Analysis via Survey and Logs
Learning Anxiety Score [^1] Anxiety was rated on a scale where lower scores are better. The difference is statistically significant ($p<.01$). 2.5 3.2
@@
Cost-Effectiveness:
EasyMED delivers comparable training at a fraction of the per-session cost of human SP, making large-scale deployment feasible.
Beyond its pedagogical benefits, EasyMED offers a substantial advantage in scalability and cost. As shown in Table 8, the per-session cost for an EasyMED interaction is approximately \$0.73, compared with \$52.95 for a human SP. This dramatic reduction in operational expense highlights the potential for EasyMED to provide high-quality clinical training at scale at only a fraction of the cost of traditional methods.
@@ 6 Conclusion
This study examines the use of large language models in SP training. We present EasyMED, a multi-agent framework that separates patient simulation, intent recognition, and evaluation into coordinated agents aligned with real clinical workflows. To enable consistent evaluation, we build the SPBench benchmark from SP–student dialogues across 14 specialties and expert-defined criteria. A four-week controlled study shows that EasyMED matches human SP training in learning outcomes and improves performance for novice learners through higher flexibility, safety, and lower cost.
Our work demonstrates the feasibility of multi-agent, LLM-based SP systems for scalable and effective clinical training. Future work will test larger cohorts, integrate multimodal interaction, and study long-term learning effects.
7 Limitations
@@ Appendix A Related Work
This section reviews research on virtual patients and intelligent tutoring, the use of large language models (LLM) in education, and automated assessment of complex, interactive skills. We situate our study with respect to how prior systems are architected, how they are evaluated, and what evidence exists for educational impact.
@@ A.1 Virtual Patients and Intelligent Tutoring
Virtual patients (VP) have long supported safe practice of diagnostic reasoning and communication in medical education cook2009virtual; berman2016role; kononowicz2015virtual. Early systems were primarily rule- or script-based, providing structured but inflexible interactions and limited behavioral realism huang2007virtual. Recent work explores LLM-driven VP to increase linguistic fluency and adaptability du2024llms; almansoori2025self; grevisse2024raspatient; steenstra2025scaffolding. However, most LLM-based VP adopt a paradigm that collapses patient simulation, dialogue control, and assessment into one model, typically focusing on the consultation/history-taking phase while leaving evaluation ad hoc and feedback coarse. Architecturally, this design makes persona stability and information disclosure emergent rather than governed; methodologically, it leaves limited support for actionable, multi-dimensional feedback. Our work differs by factorizing the interaction into coordinated agents with first-class interfaces for learner intent, disclosure policy, and inquiry coverage, and by pairing the simulator with a protocol that evaluates turn- and session-level behaviors relevant to clinical training.
@@ A.2 Large Language Models in Education
LLMs have been studied for tutoring, content generation, and role-playing across domains gan2023large; wang2024large; xu2024large; safranek2023role. In medical education, they have been used to generate patient histories and to support clinical decision making, and to simulate patient dialogues for practicing history-taking waisberg2024large; thirunavukarasu2023large; almansoori2025self; fan2023nphardeval. Persistent challenges include factual reliability, long-context consistency, and alignment with professional standards and safety constraints li2024enhancing; mahaut2024factual; gai2024achieving. Most studies report surface metrics or static outcomes (e.g., diagnosis/referral accuracy) rather than interaction competencies. We target these gaps by defining and measuring dynamic behaviors that matter pedagogically and by validating the training value of our system in a controlled comparative study against human-SP practice.
@@ A.3 Automated Assessment of Complex Skills
Automated assessment with LLMs has advanced in essay scoring and feedback zhangpersonaeval; lee2024analyzing; gusev2024pingpong and program analysis for coding education parvathy2025automated; chen2020analysis, but much of this work provides single-dimensional scores and limited interoperability with respect to process. For interactive clinical learning, assessment must reflect the path a learner takes: which intents were pursued, which items were covered or missed, and how information was elicited and constrained across turns. Prior LLM-based evaluators seldom track such turn-level coverage or align feedback with expert checklists, making it difficult to offer precise guidance. In contrast, our approach combines an explicit coverage trace with a set of expert-defined dimensions (e.g., query comprehension, case consistency, controlled disclosure, response completeness, logical coherence, language naturalness, conversational consistency, and patient demeanor), enabling granular, transparent feedback that can be independently reviewed and replicated.
@@ Appendix B Data Annotation Statement
The annotation and validation of all datasets in this study were conducted by qualified medical professionals. Three clinical experts reviewed all cases to ensure medical accuracy and representativeness. Two licensed physicians with outpatient referral experience independently annotated and verified the dialogue data according to predefined guidelines, with only consensus cases retained. In addition, three fifth-year medical students performed manual proofreading and transcription of dialogue records to correct OCR errors. For the creation of standardized patient scripts, three SP education specialists collaborated with three professional standardized patient actors to design and validate clinically authentic dialogues. All contributors participated voluntarily and provided informed consent under institutional ethical approval.
@@ Appendix C Construction of the Intent Recognition Test Dataset
The preliminary corpus was subsequently reviewed by a three-person expert panel, composed of professional medical personnel. The panel’s task was to remove any questions that were ambiguous, clinically unrealistic, or had unclear intent attribution to ensure the high quality and validity of each entry in the final dataset. After meticulous manual filtering and proofreading, we finalized a validated dataset containing 4,631 clinical questions.
@@ 31 Other Other relevant inquiries
Table 10: Performance comparison of different models on the balanced intent recognition task, reflecting realistic fluctuations. All scores are percentages (%).
Appendix D Clinical Case Data Preparation
@@
D.1 Case Source and Selection
@@
D.2 Data Processing Workflow
@@
Appendix E Outcome Measures
@@
E.1 Primary Outcome Measure: OSCE Scores
We administered OSCE tests to all students at three time points: pre-experiment, mid-experiment, and post-experiment. To avoid learning effects, the cases used in the three tests were different but were reviewed by experts to ensure consistent difficulty and assessment points. Scoring was performed independently by two blinded examiners who were unaware of the students’ group assignments, ensuring objectivity.
E.2 Secondary Outcome Measure: Subjective Questionnaire
@@
E.2.1 Part 1: Background Information

### v2

@@ 6.1 Experimental Design
Participants We recruited 20 medical undergraduate students in their fourth or fifth year. All participants completed a pre-test and received a 10-minute introduction to EasyMED. Students with scheduling conflicts or anomalous test scores were excluded (Appendix G), yielding a final cohort of 14 students (7 men, 7 women; age range 21–24, mean 23). All participants had completed core clinical coursework but had not yet taken the National Medical Licensing Examination.
Participants were ranked by pre-test scores and assigned to groups using an alternating allocation scheme. Three experienced professionals served as human standardized patients. All participants were compensated on an hourly basis in accordance with institutional ethical guidelines.
6.2 Results and Analysis
6.2.1 Evaluating Overall Improvement via OSCE
Objective Structured Clinical Examination (OSCE) is a standardized, station-based clinical skills assessment. We first confirm that the two groups were comparable prior to the intervention. As shown in Figure 5, baseline OSCE score distributions did not differ significantly between Group A (mean = 70.56) and Group B (mean = 69.84; $t(12)=0.16$, $p=0.88$), indicating similar starting levels.
Figure 5: Boxplots of baseline OSCE scores for Group A and Group B prior to the intervention. Each point represents an individual participant.
Overall Performance Across the four-week study, both groups demonstrate substantial and comparable improvements in OSCE scores. As summarized in Table 5, Group A improved by 16.89 points on average, while Group B improved by 15.36 points. Figure 4 shows consistent upward trends across individual learners in both groups.
Finding 1: These results indicate that EasyMED supports clinical skill acquisition at a level comparable to human standardized patient training.
Phase-wise Effects Most learning gains occurred during the initial training phase for both modalities. During Phase 1 (Weeks 1–2), Group A gained 15.51 points using EasyMED, while Group B gained 12.17 points using human SPs. In Phase 2 (Weeks 3–4), when groups switched modalities, additional gains were observed but at a slower rate, suggesting diminishing returns commonly seen in short-term intensive training.
Improvement by Skill Level To examine individual differences, we stratify participants into high-baseline (top three) and low-baseline (bottom four) groups based on their pre-test OSCE scores. As shown in Figure 7, low-baseline using EasyMED gained an average of 21.83 points, compared to 16.58 points with human SP. High-baseline improved less (7.10 vs. 6.30).
Finding 2: This indicates that EasyMED is particularly effective for novice learners during the early stage of training (i.e., the first two weeks).
@@ 6.2.1 Evaluating Overall Improvement via OSCE
6.2.2 Behavioral Analysis via Survey and Logs
To further contextualize the learning outcomes reported in Section 6.2.1, we analyzed students’ subjective questionnaire responses (see Appendix F) together with interaction logs collected during training.
Perceived Authenticity Students reported high perceived realism when interacting with EasyMED. On a five-point Likert scale, the simulated patient dialogue achieved a mean authenticity score of 4.6 (Table 6), indicating that the interaction was generally regarded as natural and clinically plausible.
Ratings for learning helpfulness were comparable to those of human SP training, suggesting that EasyMED is perceived not merely as a convenient substitute, but as a viable modality for practicing history-taking and clinical reasoning.
Peer Pressure Survey results indicate substantially lower learning anxiety during EasyMED sessions than during human SP interactions (mean anxiety score 0.5 vs. 3.2, $p<.01$). Students reported feeling less concerned about making mistakes and more willing to ask exploratory or repeated questions. This low-pressure environment may facilitate risk-free exploration, particularly for learners at an early stage of training.
Behavioral Evidence System logs provide objective evidence that complements these subjective reports. As shown in Figure 6, EasyMED practice sessions were distributed across a wide range of times, including evenings and weekends, whereas human SP sessions were largely confined to weekday working hours. In addition, EasyMED sessions involved more interaction on average, with a higher number of dialogue turns (54 vs. 47) and longer session durations (28:49 vs. 15:17) than human SP sessions (Table  6). Although text-based interaction may partially account for longer durations, the increased number of turns suggests more iterative questioning and sustained engagement.
@@ 6.2.2 Behavioral Analysis via Survey and Logs
Authenticity 4.6
Helpfulness 4.5 4.7
Learning Anxiety Score [^3] Anxiety was rated on a scale where lower scores are better. The difference is statistically significant ($p<.01$). 0.5 3.2
@@
Cost-Effectiveness: For cost estimation, EasyMED session costs are computed from the total token usage of a complete training interaction, whereas human SP costs are estimated by converting standard hourly compensation into a per-session cost. This results in an approximately 73-fold cost reduction compared to traditional SP training.
Finding 3: EasyMED provides a realistic, low-pressure, and accessible training environment that supports sustained and exploratory practice at a fraction of the cost of human SP training.
@@ 7 Conclusion
This study examines whether large language models can function as standardized patients for clinical skills training. We propose EasyMED, a multi-agent framework for stable, inquiry-conditioned patient simulation, and introduce SPBench, a human-grounded benchmark built from standardized patient–student dialogues. In a four-week controlled study, EasyMED achieves learning outcomes comparable to human SP training, with stronger early gains for novice learners, greater flexibility, and substantially lower cost. These results suggest that LLM-based multi-agent VSPs are a practical and scalable complement to traditional SP programs.
Limitations
@@ Appendix A Related Work
This section reviews research on virtual patients and intelligent tutoring, the use of large language models in education, and automated assessment of complex, interactive skills. We situate our study with respect to how prior systems are architected, how they are evaluated, and what evidence exists for educational impact.
@@ A.1 Virtual Patients and Intelligent Tutoring
Virtual patients (VP) have long supported safe practice of diagnostic reasoning and communication in medical education Cook and Triola (2009); Berman et al. (2016); Kononowicz et al. (2015). Early systems were primarily rule- or script-based, providing structured but inflexible interactions and limited behavioral realism Huang et al. (2007). Recent work explores LLM-driven VP to increase linguistic fluency and adaptability Du et al. (2024); Almansoori et al. (2025); Grévisse (2024); Steenstra et al. (2025). However, most LLM-based VP adopt a paradigm that collapses patient simulation, dialogue control, and assessment into one model, typically focusing on the consultation/history-taking phase while leaving evaluation ad hoc and feedback coarse. Architecturally, this design makes persona stability and information disclosure emergent rather than governed; methodologically, it leaves limited support for actionable, multi-dimensional feedback. Our work differs by factorizing the interaction into coordinated agents with first-class interfaces for learner intent, disclosure policy, and inquiry coverage, and by pairing the simulator with a protocol that evaluates turn- and session-level behaviors relevant to clinical training.
@@ A.2 Large Language Models in Education
LLMs have been studied for tutoring, content generation, and role-playing across domains Gan et al. (2023); Wang et al. (2024); Xu et al. (2024); Safranek et al. (2023). In medical education, they have been used to generate patient histories and to support clinical decision making, and to simulate patient dialogues for practicing history-taking Waisberg et al. (2024); Thirunavukarasu et al. (2023); Almansoori et al. (2025); Fan et al. (2023). Persistent challenges include factual reliability, long-context consistency, and alignment with professional standards and safety constraints Li et al. (2024a); Mahaut et al. (2024); Gai et al. (2024). Most studies report surface metrics or static outcomes (e.g., diagnosis/referral accuracy) rather than interaction competencies. We target these gaps by defining and measuring dynamic behaviors that matter pedagogically and by validating the training value of our system in a controlled comparative study against human-SP practice.
@@ A.3 Automated Assessment of Complex Skills
Automated assessment with LLMs has advanced in essay scoring and feedback Zhang et al. ; Lee et al. (2024); Gusev (2024) and program analysis for coding education Parvathy et al. (2025); Chen et al. (2020), but much of this work provides single-dimensional scores and limited interoperability with respect to process. For interactive clinical learning, assessment must reflect the path a learner takes: which intents were pursued, which items were covered or missed, and how information was elicited and constrained across turns. Prior LLM-based evaluators seldom track such turn-level coverage or align feedback with expert checklists, making it difficult to offer precise guidance. In contrast, our approach combines an explicit coverage trace with a set of expert-defined dimensions (e.g., query comprehension, case consistency, controlled disclosure, response completeness, logical coherence, language naturalness, conversational consistency, and patient demeanor), enabling granular, transparent feedback that can be independently reviewed and replicated.
@@ Appendix B Data Annotation Statement
To ensure medical accuracy, pedagogical validity, and ethical compliance, we assembled a multidisciplinary team composed of clinical experts, licensed physicians, medical students, and standardized patient (SP) professionals. The specific roles and contributions were distributed as follows:
Clinical Expert Panel
A panel of three clinical experts, consisting of two senior physicians with eight years of clinical experience and one attending physician with five years of experience, was responsible for the high-level design and validation of the study. Their duties included defining the eight expert evaluation criteria for SPBench, validating the intent recognition dataset, and overseeing the selection of clinical cases from authoritative training sources.
Data Annotation and Blind Review
To ensure objectivity and inter-rater reliability, specific annotation tasks were conducted by two independent licensed physicians (with three and five years of clinical experience, respectively) who were blinded to the model outputs and student groupings. Their specific tasks included:
  • Case Quality Control: Checking every processed case to ensure anonymity, realism, and teaching utility.
  • Benchmarking: Conducting a blind review of 86 randomly selected samples to validate the automated GPT-4o evaluation scores.
  • OSCE Scoring: Independently scoring the pre-, mid-, and post-experiment OSCE tests for all student participants.
  • Evaluation Agent Validation: Assessing the clinical appropriateness and validity of the guidance generated by the Evaluation Agent across 30 distinct practice sessions (as detailed in Appendix  D).
Data Processing and Annotation Support
Three senior medical students (5th-year undergraduates) were recruited for data preparation tasks.
  • Digitization: They performed manual proofreading and correction of OCR-scanned text from the source books to fix typos and punctuation errors.
  • Auxiliary Agent Validation: They participated in the construction of the Intent Recognition Test Dataset (Appendix  C). This involved reviewing and filtering the preliminary corpus of clinical questions generated by GPT-4o to remove ambiguous or unrealistic entries, ensuring the dataset’s quality for testing the Auxiliary Agent.
Standardized Patient Script and Simulation
The creation of high-fidelity scripts and human SP performance involved a collaborative team of three SP education instructors (specializing in 5th-year medical student training) and three professional SPs (with two years of acting experience). This team designed the patient history, symptoms, and emotional cues. The same three experienced professionals served as the human SPs during the four-week comparative user study.
Ethical Compliance
All contributors, including students, actors, and experts, participated with informed consent. They were compensated for their time adhering to standard hourly wage practices.
@@ Appendix C Construction of the Intent Recognition Test Dataset
The preliminary corpus was subsequently reviewed by a three medical student panel, composed of professional medical personnel. The panel’s task was to remove any questions that were ambiguous, clinically unrealistic, or had unclear intent attribution to ensure the high quality and validity of each entry in the final dataset. After meticulous manual filtering and proofreading, we finalized a validated dataset containing 4,631 clinical questions.
@@ 31 Other Other relevant inquiries
Table 8: Performance comparison of different models on the intent recognition task. All scores are reported as percentages (%).
Appendix D Validation of the Evaluation Agent Guidance
To ensure that the post-session guidance provided by the Feedback Agent (e.g., highlighting missed or superfluous inquiries) is clinically sound and pedagogically appropriate, we conducted an independent validation study.
Study Design and Methodology
We sampled 60 complete practice sessions from the user study. For each session, the specific feedback generated by the agent was extracted and anonymized. Two independent clinical experts, blinded to the source of the generation, rated the appropriateness of each feedback item on a 5-point Likert scale (1=misleading, 5=highly appropriate). We defined two primary evaluation metrics: Accuracy, calculated as the percentage of feedback items receiving a score of $\geq 4$ from both experts; and Inter-rater Agreement, measured using Cohen’s $\kappa$.
Results and Discussion
The validation results are summarized in Table 9. Across the 60 evaluations, the Feedback Agent demonstrated high reliability, achieving an accuracy of 87% with substantial agreement between experts ($\kappa=0.76$). Qualitative error analysis revealed that most disagreements arose in borderline cases where the clinical necessity of a specific inquiry was debatable. These findings confirm that the agent provides generally reliable guidance. Future iterations could incorporate confidence scores to allow experts to flag ambiguous feedback for refinement.
Table 9: Validation results of the Feedback Agent’s guidance across 60 sessions. Accuracy is defined as the percentage of items rated $\geq 4$ by both clinical experts.
Metric Value Description
Sample Size 60 Total number of sessions evaluated
Accuracy 87% Proportion of feedback rated as appropriate
Inter-rater Agreement 0.76 Cohen’s $\kappa$ indicating expert consensus
Appendix E Clinical Case Data Preparation
@@
E.1 Case Source and Selection
@@
E.2 Data Processing Workflow
@@
Appendix F Outcome Measures
@@
F.1 Primary Outcome Measure: OSCE Scores
We administered OSCE tests to all students at three time points: pre-experiment, mid-experiment, and post-experiment. To avoid learning effects, the cases used in the three tests were different but were reviewed by experts to ensure consistent difficulty and assessment points. Scoring was performed independently by two blinded examiners who were unaware of the students’ group assignments, ensuring objectivity. All participants provided written informed consent prior to participation.
F.2 Secondary Outcome Measure: Subjective Questionnaire
@@
F.2.1 Part 1: Background Information